OpenStudy (anonymous):

Solve 2x2 - 8x = -7.

OpenStudy (anonymous):

4 - 8x = -7 -8x = -7 - 4 -8x = -11 8x = 11 x = 11/8 x = 1 3/8 or 1.375

OpenStudy (anonymous):

these are my choices

OpenStudy (mertsj):

1. Add 7 to both sides. 2. Plug into the quadratic formula 3. a = 2, b = -8, c = 7

OpenStudy (anonymous):

Is the 2x squared (see below) or does it say 2 times 2?\[2x ^{2}\]

OpenStudy (anonymous):

its 2x squared (:

OpenStudy (anonymous):

ahhhhh... 2x^2, very different! LOL hold on

OpenStudy (anonymous):

hahah alright (:

OpenStudy (mertsj):

@olivia0311 What did you get when you plugged into the formula?

OpenStudy (anonymous):

\[x=8\pm \sqrt{8} \over 4 \]

OpenStudy (mertsj):

Yes. \[\frac{8\pm \sqrt{8}}{4}\]

OpenStudy (mertsj):

So divide out the 4

OpenStudy (mertsj):

And simplify radical 8

OpenStudy (anonymous):

\[-2\pm 2 \sqrt{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[\frac{8\pm \sqrt{8}}{4}=\frac{8\pm2\sqrt{2}}{4}=\frac{4\pm \sqrt{2}}{2}=2\pm\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

oh okay, I see what I did wrong, thank you so much